OpenStudy (anonymous):

A movie theatre has a capacity of 41 seats. When there are d spectators in the theatre, there are d+13 empty seats. With t spectators, there are d+ 19 empty seats. What is the values of t?

OpenStudy (anonymous):

41 seats, separated in two categories : empty and used. 1) d spectators 41=(d) + (d+13) d=14 2) t spectators 41=(t)+(d+19) t=8

OpenStudy (anonymous):

ty so much

OpenStudy (anonymous):

no problem!